Use the References to access important values if needed for this question. The following initial rate data are for the reaction of ammonium ion with nitrite ion in aqueous solution NH, NO 2 Ho 2req 780 780 1.56 1.56 6.56 102 1.54 10-5 07 105 3.07 10-5 6.13 10 0.131 56x10 131 ts 2req Complete the rate law for this reaction in the box below Use the form kIAl(B, where 1 is understood for m or n and concentrations taken to the zero power do not appear: Dont enter 1 for m orn pts From these data, the rate constant is M11 Submit Aswe Retry Entire Groug 9 more

media%2Fad5%2Fad535d48-ea23-48ab-b194-9b

media%2Fede%2Fede753c6-a653-424b-8378-a7

help it this 3 questions please
Answer #1

1.

Do, (1) / (2)

1.54 * 10-5 / (3.07 * 10-5) = (6.56 * 10-2 / 0.131)n

0.502 = (0.501)n

n = 1

Therefore, order with respect to NO2- = 1

Now do ((1) / (3)

1.54 * 10-5 / (3.07 * 10-5) = (0.780 / 1.56)m

0.502 = (0.5)m

So, m = 1

Therefore, order with respect to NH4+ = 1

Over order of reaction = 1 + 1 = 2

Rate law expression can be written as,

Rate = k[NH4+][NO2-]

From (1)

1.54 * 10-5 = k(0.780)(6.56 * 10-2)

k = 3.01 * 10-4 M-1s-1